It took Andy 9 hours to drive to a rock concert. On the way home, he was able to increase his average speed by 14 mph and make the return drive in only 7 hours. What is his average speed for the drive home
Answer by ikleyn(52786)

.
It took Andy 9 hours to drive to a rock concert. On the way home, he was able to increase his average speed by 14 mph and
make the return drive in only 7 hours. What is his average speed for the drive home?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Your guiding equation is 

9*u = 7*(u+14),

where u is his speed on the way to the rock concert, in mph.  (Each side of the equation is the distance between home and concert).

The solution is u = 49 mph.

So, the speed on the way home is 49+14 = 63 mph.